A scooterist is racing at a speed of 72 km per hour. If the radius of the wheel is 20 cm, find the angular speed of the wheels.

Grade:10

2 Answers

MURALI KRISHNA YADAV
11 years ago

72kmph=20m/s

20cm=0.2m;

v=rw

w=v/r

w=20/0.2

=100rad/sec is the answer..just a basic question think for it...... u can do miracles............

kumar sanu
6 years ago
We know that 72km/h=20m/s20cm=0.2mV=R/WW=V/RW=20/0.2W=100rad/s this is hard but not for those who are a good student